OpenStudy (anonymous):

1.A red number cube and a blue number cube are rolled. If all numbers are equally likely, what is the probability that the sum of the numbers is 6? Enter your answer in simplified fraction form; example: 3/20 2.A red number cube and a blue number cube are rolled. If all numbers are equally likely, what is the probability that the difference of the numbers is 6? Enter a numerical answer.

OpenStudy (mathmate):

Do you know the total number of outcomes?

OpenStudy (mathmate):

... when you throw two different number cubes?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

P(sum is 6)=3/36=1/12?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@myininaya @ganeshie8

OpenStudy (anonymous):

@abb0t @aaronq @zepdrix

OpenStudy (anonymous):

actually there are 5 outcomes... 1,5 5,1 2,4 4,2 3,3

OpenStudy (anonymous):

you are the most help ever, thank you so much!

OpenStudy (anonymous):

your welcome anytime

OpenStudy (mathmate):

I hope you get the second part without much trouble. :)